# On the Use of Agentic Coding: An Empirical Study of Pull Requests on GitHub

### v1

@@ 5.1. Implications for researchers
Create PR-centric benchmarks that reward alignment across code, documentation, and review resolution. Current benchmarks such as SWE-bench (Jimenez et al., 2024) overlook documentation drift, style adherence, and reviewer negotiation, all of which accounted for 48.5% of Agentic-PR revisions (Section 4.4). This gap between benchmark evaluation and real-world requirements represents a fundamental misalignment in how we assess coding agent capabilities. While existing benchmarks focus predominantly on functional correctness, whether the generated code passes test cases, they fail to capture nearly half of the actual work required to integrate AI-generated code into production systems. We suggest building datasets that incorporate time-to-merge, rework size, and the alignment between documentation and implementation. Such benchmarks can be applied during both training and evaluation of coding agents to promote improvements beyond functional correctness.
@@ 5.2. Implications for developers
Split large, multi-purpose PRs into smaller, easy-to-review submissions. Over 27% of Agentic-PRs combined multiple tasks (Section 4.1), and “too large” was among the top three rejection reasons (Section 4.2). Large PRs increase reviewer cognitive load and raise the risk of rejection, as previous studies (Weißgerber et al., 2008) have warned developers for decades. This challenge remains relevant even for agents. In fact, RQ1 shows agents generate comprehensive solutions that attempt to address multiple issues simultaneously. The tendency of AI coding agents to produce extensive changes when given broad directives mirrors the problematic patterns seen in human development, but with potentially greater scale and complexity. When assigning a task to a coding agent, developers should adhere to the principle of providing a small, self-contained task that can be addressed in a single PR, or structure broader tasks to produce a sequence of smaller, follow-up PRs for wide-reaching changes.
@@ 5.3. Implications for coding agent builders
Coding agent builders should integrate additional tools to improve adherence to project conventions. The high rates of style (22.122.1%) and refactoring (25.725.7%) revisions indicate that generic models do not fully capture local conventions (Section 4.4). This also implies that creating a persistent friction between agent capabilities and team expectations. This misalignment stems from agents being trained on diverse codebases with conflicting conventions, making them unable to infer project-specific patterns from limited context. Using the Model Context Protocol (MCP) servers (Hasan et al., 2025), builders can connect agents to linters (e.g., pylint) and static analysis tools to enforce formatting and structural rules before submission. Similarly, build-checking and test-coverage tools could be integrated to catch common revision triggers in advance.
@@ 6.1. Human-AI collaboration in software engineering
Recent research has emphasized the increasing collaboration between human developers and AI-powered technologies in common software engineering tasks (Lambiase et al., 2025), such as requirements documentation (Rahman et al., 2024), code search (Liu et al., 2024b), and code generation (Dakhel et al., 2023). Hassan et al. (Hassan et al., 2024) characterized this collaborative relationship within the emerging paradigm of Software Engineering 3.0, envisioning AI teammates that enhance developer productivity and reduce the cognitive load on human developers. However, despite these technical advances, human oversight remains crucial. AI-powered tools have demonstrated potential for improving productivity, but human-in-the-loop approaches are still required to maintain high-quality results (Coutinho et al., 2024; Mock et al., 2024). Hence, establishing clear guidelines and best practices for managing human-AI interactions remains an essential step for industrial organizations to integrate these technologies (Santos et al., 2024).
In addition, multi-agent frameworks such as LLM-Agent, FlowGen and FightFire assign specialised roles (planner, coder, tester) to cooperating LLMs, emulating Agile practices and raising robustness (He et al., 2025; Lin et al., 2025; Yu et al., 2024). A recent systematic review consolidates prompt-engineering techniques and taxonomies of human-AI collaboration scenarios in software engineering (Hou et al., 2024).
However, despite these technical achievements, significant social and human challenges exist in AI-assisted development. Piorkowski et al. (Piorkowski et al., 2021) and Nahar et al. (Nahar et al., 2025) uncover persistent knowledge gaps inside multi-disciplinary AI teams, reinforcing the call for explicit interaction guidelines. Furthermore, Adam et al. (Alami et al., 2025) found that LLM-assisted reviews can disrupt collective accountability mechanisms inherent in traditional peer reviews, as developers value intrinsic factors such as professional integrity and reputation that cannot be maintained when interacting with AI systems. These studies highlight the complex socio-technical challenges in human-AI collaboration.
@@ 6.2. AI-assisted issue reports and PRs
Previous studies have explored AI-assisted activities related to issue reports (Bo et al., 2024; Kang et al., 2023) and PRs (Chouchen et al., 2024; Watanabe et al., 2024). However, no prior research has investigated the use of agentic coding tools, such as OpenAI’s Codex [^47] https://openai.com/index/introducing-codex and Claude Code, in real-world development contexts. Instead, current research primarily focuses on LLM-based chatbots like ChatGPT. For example, Chouchen et al. (Chouchen et al., 2024) reported that ChatGPT is frequently used in review-intensive PRs to reduce developer effort. Despite this positive impact, 30.7% of AI-generated suggestions during code reviews were still met with skepticism or disagreement (Watanabe et al., 2024). Our study fills this gap by tracing 567 autonomous PRs created by Claude Code across 157 open-source projects, reporting acceptance rates, revision effort, and rejection rationales.
Researchers have proposed various approaches for directly involving LLMs in activities related to issue reports and PRs. For example, Kang et al. (Kang et al., 2023) used LLMs to generate test cases from issue reports for easier bug reproduction, Zhang et al. (Irsan et al., 2022) trained LLMs to automatically generate PR titles, and Bo et al. (Bo et al., 2024) leveraged ChatGPT to improve the clarity of issue reports for helping subsequent tasks such as bug fixing.
Complementary work investigates AI agents that clarify ambiguous requirements and collaboratively co-evolve PR descriptions; for instance, AgileGen aligns end-user perspectives with acceptance criteria (Zhang et al., 2025), while large-scale analyses of LLMs on code-change tasks reveal their potential for automated review of PR diffs (Fan et al., 2025). Multi-agent benchmarks also examine ChatGPT’s ability to self-verify its generated code through test-report generation, reducing reviewer burden (Yu et al., 2024).
Early work by Tufano et al. (Tufano et al., 2021) pioneered DL models that propose code changes before review submission and apply reviewer suggestions after submission. Guo et al. report that GPT-4 can refine 66% of files given review comments (Guo et al., 2024), and later boost intent-extraction accuracy to 79% in their ICSE ’25 study (Guo et al., 2025). To improve developer confidence, Di and Zhang (Di and Zhang, 2025) interleave inline comments with LLM edits, cutting task time by 16.7 percent.
@@ 6.3. Empirical studies of automatic code generation
Researchers have studied various aspects of AI-generated code, including correctness (Billah et al., 2024), efficiency (Niu et al., 2024), and security (Liu et al., 2024a). For example, Billah et al. (Billah et al., 2024) evaluated the correctness (pass rate) of LLMs in solving programming challenges on competitive programming platforms such as LeetCode and Codeforces. They reported a decrease in the pass rate in live Codeforces contests compared to archived problems. However, code correctness alone does not necessarily imply efficiency or optimality as code complexity increases, and extra effort is required to support AI models in generating efficient code (Niu et al., 2024). From a security perspective, Liu et al. (Liu et al., 2024a) identified security vulnerabilities in AI-generated code and highlighted potential security risks that must be mitigated before deployment in production environments. Beyond generating code at the granularity of individual functions or files, SWE-bench (Jimenez et al., 2024) was introduced to evaluate the performance of AI-powered code generation tools in resolving real-world GitHub issues.
Subsequent work has shown that ChatGPT often produces buggy or low-quality code that benefits from iterative repair (Liu et al., 2024a), while self-collaboration frameworks with multiple LLM agents significantly improve correctness and reduce human intervention (Dong et al., 2024). Empirical studies highlight the challenges of domain-specific coding (e.g., machine-learning pipelines) and demonstrate that knowledge-prompting strategies boost performance in such settings (Gu et al., 2025; Shin et al., 2024). Recent surveys map out future research directions for automatic programming pipelines, including integrated quality-assurance stages (Lyu et al., 2025), and bias-testing frameworks reveal that LLMs can inherit societal biases, motivating prompt-engineering mitigation strategies (Huang et al., 2025).
Wang et al. (Wang et al., 2025) catalogue 557 distinct error patterns that six LLMs make on HumanEval. Niu et al. benchmark 19 pre-trained models across 13 tasks, confirming that functional correctness rarely correlates with efficiency (Niu et al., 2023). O’Brien et al. (O’Brien et al., 2024) show Copilot frequently injects TODO-style self-admitted technical debt. For repair, Xia et al. (Xia et al., 2023) and Bouzenia et al. (Bouzenia et al., 2025) re-cast bug fixing as an autonomous planning problem executed by an LLM agent.
@@ 7.1. Threats to internal validity
A potential concern in our study is the ambiguity related to developers’ behaviors when using Claude Code. Developers may modify the generated code before committing it. Conversely, even when developers use Claude Code to generate code, they may directly push the changes without requesting the tool to perform the commit itself. To mitigate this threat, we explicitly included only PRs clearly marked as “Generated with Claude Code.” This explicit labeling provides a strong signal of Claude Code’s involvement, reducing potential ambiguity about code provenance.
@@ 7.2. Threats to construct validity
Claude Code implements agentic coding and can leverage the Model Context Protocol (MCP), which enables agents to integrate with external tools and data sources. However, this study does not isolate the specific contributions of agentic capabilities and MCP integration to the observed outcomes. Disentangling their individual impacts remains an important direction for future research.
@@ 7.3. Threats to external validity
This study examines only 567 PRs that explicitly used Claude Code, which represents the full set we were able to retrieve. One reason for this limited sample size is that Claude Code was launched only recently, in February 2025. To validate our findings, replication studies with larger datasets collected over a longer time frame will be necessary. Furthermore, our analysis focused exclusively on Claude Code, an agentic coding tool whose usage could be reliably identified in PRs. As a result, our findings may not generalize to other agentic coding tools. Still, we believe that the findings here may be indicative of broader trends, given that these functionalities and performance are similar.
@@ Acknowledgements.
We gratefully acknowledge the financial support of JSPS KAKENHI grants (JP24K02921, JP25K21359), as well as JST PRESTO grant (JPMJPR22P3), ASPIRE grant (JPMJAP2415), and AIP Accelerated Program (JPMJCR25U7).

### v2

@@ 5.1. Implications for researchers
Create PR-centric benchmarks that reward alignment across code, documentation, and review resolution. Current benchmarks such as SWE-bench (DBLP:conf/iclr/JimenezYWYPPN24) overlook documentation drift, style adherence, and reviewer negotiation, all of which accounted for 48.5% of Agentic-PR revisions (Section 4.4). This gap between benchmark evaluation and real-world requirements represents a fundamental misalignment in how we assess coding agent capabilities. While existing benchmarks focus predominantly on functional correctness, whether the generated code passes test cases, they fail to capture nearly half of the actual work required to integrate AI-generated code into production systems. We suggest building datasets that incorporate time-to-merge, rework size, and the alignment between documentation and implementation. Such benchmarks can be applied during both training and evaluation of coding agents to promote improvements beyond functional correctness.
@@ 5.2. Implications for developers
Split large, multi-purpose PRs into smaller, easy-to-review submissions. Over 27% of Agentic-PRs combined multiple tasks (Section 4.1), and “too large” was among the top three rejection reasons (Section 4.2). Large PRs increase reviewer cognitive load and raise the risk of rejection, as previous studies (DBLP:conf/msr/WeissgerberND08) have warned developers for decades. This challenge remains relevant even for agents. In fact, RQ1 shows agents generate comprehensive solutions that attempt to address multiple issues simultaneously. The tendency of AI coding agents to produce extensive changes when given broad directives mirrors the problematic patterns seen in human development, but with potentially greater scale and complexity. When assigning a task to a coding agent, developers should adhere to the principle of providing a small, self-contained task that can be addressed in a single PR, or structure broader tasks to produce a sequence of smaller, follow-up PRs for wide-reaching changes.
@@ 5.3. Implications for coding agent builders
Coding agent builders should integrate additional tools to improve adherence to project conventions. The high rates of style (22.122.1%) and refactoring (25.725.7%) revisions indicate that generic models do not fully capture local conventions (Section 4.4). This also implies that creating a persistent friction between agent capabilities and team expectations. This misalignment stems from agents being trained on diverse codebases with conflicting conventions, making them unable to infer project-specific patterns from limited context. Using the Model Context Protocol (MCP) servers (hasan2025mcp), builders can connect agents to linters (e.g., pylint) and static analysis tools to enforce formatting and structural rules before submission. Similarly, build-checking and test-coverage tools could be integrated to catch common revision triggers in advance.
@@ 6.1. Human-AI collaboration in software engineering
Recent research has emphasized the increasing collaboration between human developers and AI-powered technologies in common software engineering tasks (DBLP:journals/csur/LambiaseCPF25), such as requirements documentation (DBLP:conf/icsm/RahmanZMRRS24), code search (DBLP:conf/sigsoft/LiuZ0WH024), and code generation (DBLP:journals/jss/DakhelMNKDJ23). Hassan et al. (DBLP:journals/corr/abs-2410-06107) characterized this collaborative relationship within the emerging paradigm of Software Engineering 3.0, envisioning AI teammates that enhance developer productivity and reduce the cognitive load on human developers. However, despite these technical advances, human oversight remains crucial. AI-powered tools have demonstrated potential for improving productivity, but human-in-the-loop approaches are still required to maintain high-quality results (DBLP:conf/aiware/CoutinhoMSDFS24; DBLP:conf/xpu/MockMR24). Hence, establishing clear guidelines and best practices for managing human-AI interactions remains an essential step for industrial organizations to integrate these technologies (DBLP:conf/icst/SantosSMS24).
In addition, multi-agent frameworks such as LLM-Agent, FlowGen and FightFire assign specialised roles (planner, coder, tester) to cooperating LLMs, emulating Agile practices and raising robustness (DBLP:journals/tosem/HeTL25; DBLP:conf/icse/LinKC25; DBLP:journals/tse/YuLHKLX24). A recent systematic review consolidates prompt-engineering techniques and taxonomies of human-AI collaboration scenarios in software engineering (DBLP:journals/tosem/HouZLYWLLLGW24).
However, despite these technical achievements, significant social and human challenges exist in AI-assisted development. Piorkowski et al. (DBLP:journals/pacmhci/PiorkowskiPWWMP21) and Nahar et al. (DBLP:conf/icse/NaharZLZK25) uncover persistent knowledge gaps inside multi-disciplinary AI teams, reinforcing the call for explicit interaction guidelines. Furthermore, Adam et al. (10.1145/3721127) found that LLM-assisted reviews can disrupt collective accountability mechanisms inherent in traditional peer reviews, as developers value intrinsic factors such as professional integrity and reputation that cannot be maintained when interacting with AI systems. These studies highlight the complex socio-technical challenges in human-AI collaboration.
@@ 6.2. AI-assisted issue reports and PRs
Previous studies have explored AI-assisted activities related to issue reports (DBLP:conf/kbse/BoJ000024; DBLP:conf/icse/KangYY23) and PRs (DBLP:conf/msr/ChouchenBBOAM24; DBLP:conf/ease/WatanabeK0HYI24). However, no prior research has investigated the use of agentic coding tools, such as OpenAI’s Codex [^47] https://openai.com/index/introducing-codex and Claude Code, in real-world development contexts. Instead, current research primarily focuses on LLM-based chatbots like ChatGPT. For example, Chouchen et al. (DBLP:conf/msr/ChouchenBBOAM24) reported that ChatGPT is frequently used in review-intensive PRs to reduce developer effort. Despite this positive impact, 30.7% of AI-generated suggestions during code reviews were still met with skepticism or disagreement (DBLP:conf/ease/WatanabeK0HYI24). Our study fills this gap by tracing 567 autonomous PRs created by Claude Code across 157 open-source projects, reporting acceptance rates, revision effort, and rejection rationales.
Researchers have proposed various approaches for directly involving LLMs in activities related to issue reports and PRs. For example, Kang et al. (DBLP:conf/icse/KangYY23) used LLMs to generate test cases from issue reports for easier bug reproduction, Zhang et al. (DBLP:conf/icsm/IrsanZT0J22) trained LLMs to automatically generate PR titles, and Bo et al. (DBLP:conf/kbse/BoJ000024) leveraged ChatGPT to improve the clarity of issue reports for helping subsequent tasks such as bug fixing.
Complementary work investigates AI agents that clarify ambiguous requirements and collaboratively co-evolve PR descriptions; for instance, AgileGen aligns end-user perspectives with acceptance criteria (DBLP:journals/tosem/ZhangXGXCZZFZ25), while large-scale analyses of LLMs on code-change tasks reveal their potential for automated review of PR diffs (DBLP:journals/tosem/FanLLLXL25). Multi-agent benchmarks also examine ChatGPT’s ability to self-verify its generated code through test-report generation, reducing reviewer burden (DBLP:journals/tse/YuLHKLX24).
Early work by Tufano et al. (DBLP:conf/icse/TufanoPTPB21) pioneered DL models that propose code changes before review submission and apply reviewer suggestions after submission. Guo et al. report that GPT-4 can refine 66% of files given review comments (DBLP:conf/icse/GuoCXLL0024), and later boost intent-extraction accuracy to 79% in their ICSE ’25 study (DBLP:conf/icse/GuoXL00B25). To improve developer confidence, Di and Zhang (DBLP:conf/icse/Di025) interleave inline comments with LLM edits, cutting task time by 16.7 percent.
@@ 6.3. Empirical studies of automatic code generation
Researchers have studied various aspects of AI-generated code, including correctness (DBLP:conf/esem/BillahRCR24), efficiency (DBLP:conf/forge/NiuZL0024), and security (DBLP:journals/tse/LiuTLZZ24). For example, Billah et al. (DBLP:conf/esem/BillahRCR24) evaluated the correctness (pass rate) of LLMs in solving programming challenges on competitive programming platforms such as LeetCode and Codeforces. They reported a decrease in the pass rate in live Codeforces contests compared to archived problems. However, code correctness alone does not necessarily imply efficiency or optimality as code complexity increases, and extra effort is required to support AI models in generating efficient code (DBLP:conf/forge/NiuZL0024). From a security perspective, Liu et al. (DBLP:journals/tse/LiuTLZZ24) identified security vulnerabilities in AI-generated code and highlighted potential security risks that must be mitigated before deployment in production environments. Beyond generating code at the granularity of individual functions or files, SWE-bench (DBLP:conf/iclr/JimenezYWYPPN24) was introduced to evaluate the performance of AI-powered code generation tools in resolving real-world GitHub issues.
Subsequent work has shown that ChatGPT often produces buggy or low-quality code that benefits from iterative repair (DBLP:journals/tse/LiuTLZZ24), while self-collaboration frameworks with multiple LLM agents significantly improve correctness and reduce human intervention (DBLP:journals/tosem/DongJJL24). Empirical studies highlight the challenges of domain-specific coding (e.g., machine-learning pipelines) and demonstrate that knowledge-prompting strategies boost performance in such settings (DBLP:journals/tosem/GuCLHZWWXW25; DBLP:journals/tosem/ShinWWSW24). Recent surveys map out future research directions for automatic programming pipelines, including integrated quality-assurance stages (DBLP:journals/tosem/LyuRRTT25), and bias-testing frameworks reveal that LLMs can inherit societal biases, motivating prompt-engineering mitigation strategies (10.1145/3724117).
Wang et al. (DBLP:conf/icse/WangZSHCMZ25) catalogue 557 distinct error patterns that six LLMs make on HumanEval. Niu et al. benchmark 19 pre-trained models across 13 tasks, confirming that functional correctness rarely correlates with efficiency (DBLP:conf/icse/NiuLNCGL23). O’Brien et al. (DBLP:conf/icse/OBrienBIASR24) show Copilot frequently injects TODO-style self-admitted technical debt. For repair, Xia et al. (DBLP:conf/icse/XiaWZ23) and Bouzenia et al. (DBLP:conf/icse/BouzeniaDP25) re-cast bug fixing as an autonomous planning problem executed by an LLM agent.
@@ 7.1. Threats to internal validity
A potential concern in our study is the ambiguity related to developers’ behaviors when using Claude Code. Developers may modify the generated code before committing it. Conversely, even when developers use Claude Code to generate code, they may directly push the changes without requesting the tool to perform the commit itself.
To mitigate this threat, we explicitly included only PRs clearly marked as ”Generated with Claude Code.” While this explicit labeling provides a strong signal of Claude Code’s involvement and reduces potential ambiguity about code provenance, we acknowledge that it implies a binary classification that may overlook hybrid workflows. Regarding the Human baseline, the presence of undisclosed AI-generated code would likely dilute the differences between the two groups. Therefore, the significant differences observed in our study likely represent conservative estimates of the distinct characteristics of agentic workflows. Furthermore, relying on brittle text markers (like “Generated with Claude Code”) is insufficient for analyzing these hybrid workflows. Future research requires more fine-grained provenance tracking mechanisms to accurately separate agent-generated code from human modifications.
Another threat arises from the temporal discrepancy in data collection, which could introduce bias due to project evolution or seasonal factors. Regarding the baseline construction, we expanded the sampling window for Human-PRs backward until the sample size matched that of Agentic-PRs, resulting in a collection period spanning approximately 10 months (back to April 2024). Our analysis of the temporal distribution shows that 42.0% (238/567 PRs) of Human-PRs were created within the exact same timeframe as Agentic-PRs (February 24 – April 30, 2025). To address potential concerns about topical or seasonal drift, we conducted a sensitivity analysis comparing this time-aligned subset against the full baseline. We found that the distribution of task types in this subset remained consistent with Table 1. For example, “fix” and “feat” remained the dominant categories (27.0% and 24.7%, respectively), and the ratio of “test” (9.0%) remained significantly lower than that of Agentic-PRs (18.8%). Furthermore, performance metrics relevant to Table 2 showed consistency; the subset achieved a comparable acceptance rate (87.0% vs. 91.0% for the full set) and median merge time (0.7h vs. 1.0h). These results suggest that our findings are robust and not significantly biased by the extended sampling window.
@@ 7.2. Threats to construct validity
Claude Code implements agentic coding and can leverage the Model Context Protocol (MCP), which enables agents to integrate with external tools and data sources. However, this study does not isolate the specific contributions of agentic capabilities and MCP integration to the observed outcomes. Disentangling their individual impacts remains an important direction for future research. Furthermore, our dataset excludes only open PRs to focus on contributions with definitive outcomes (merged or rejected). This exclusion creates a survival bias, as complex or controversial PRs that remain open for extended periods are not captured. Consequently, the reported acceptance rates and merge times reflect only completed workflows and may underestimate the latency of ongoing, undecided contributions.
Finally, our reliance on PR acceptance as a primary success metric does not account for long-term code quality or maintainability. A merged PR may still introduce latent bugs or negatively impact the codebase over time (e.g., causing regressions or necessitating subsequent reverts). Since our study focuses on the immediate outcome of the review process, we lack longitudinal data to assess these downstream effects. Future research should investigate the long-term stability and maintainability of agent-generated code through longitudinal analysis.
@@ 7.3. Threats to external validity
This study examines only 567 PRs that explicitly used Claude Code, which represents the full set we were able to retrieve. One reason for this limited sample size is that Claude Code was launched only recently, in February 2025. Consequently, some observed developer behaviors may reflect an “enthusiastic phase” characterized by experimentation (e.g., PR [^48] https://github.com/thefrontside/effection/pull/992 explicitly noting “I wanted to try Claude Code”). These patterns may differ from steady-state norms that emerge once the tool is fully integrated into standard workflows. Future work should expand this time window to determine if these patterns persist as the technology matures.
To validate our findings, replication studies with larger datasets collected over a longer time frame will be necessary. Finally, our analysis focused exclusively on Claude Code, an agentic coding tool whose usage could be reliably identified in PRs. As a result, our findings may not generalize to other agentic coding tools. We view this as a critical avenue for future research and encourage the community to investigate how different agentic coding tools (e.g., GitHub Copilot) influence PR characteristics and acceptance rates.
@@ Acknowledgements.
We gratefully acknowledge the financial support of JSPS KAKENHI grants (JP24K02921, JP25K21359), as well as JST PRESTO grant (JPMJPR22P3), ASPIRE grant (JPMJAP2415), and AIP Accelerated Program (JPMJCR25U7). We also acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).